OpenStudy (anonymous):

Express the complex number in trigonometric form. 4i

OpenStudy (anonymous):

HEY NICE PIC

OpenStudy (anonymous):

NEED SOME HELP

OpenStudy (anonymous):

\(4(cos(\pi/2)+i\sin(\pi/2))\)